Circular No. 8454

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004gt, 2004gu, 2004gv The discoveries of three appparent supernovae have been reported:

2004gt by L. A. G. Monard (cf. IAUC 8430), 2004gu by R. Quimby, C. Gerardy, P. Hoeflich, and J. C. Wheeler (cf. IAUC 8446; note that Gerardy is at Imperial College), and 2004gv by Y.-t. Chen (cf. IAUC 8420).

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004gt Dec. 12.076 12^{h}01^{m}50.^{s}37 −18^{o}52′12′′.7 14.9 34′′ W, 10′′ S
2004gu Dec. 13.50 12 46 24.72 +11 56 56.1 17.5 1′′.1 W, 2′′.1 S
2004gv Dec. 13.669 2 13 37.42 − 0 43 05.8 17.6 13′′.8 W, 4′′.0 S

Additional unfiltered CCD magnitudes by the respective discoverers: SN 2004gt in NGC 4038, May 11.250 UT, [15.7 (dense region); Dec. 16.015, 14.6. SN 2004gu in FGC 175A (cf. Karachentsev et al. 1993, A.N. 314, 97), June 15.17, [17.7; Dec. 15.49, 17.5. SN 2004gv in NGC 856, Nov. 28.742, [18.6; Dec. 14.569, 17.4. Monard’s offset for 2004gt above is with respect to the more northern of the double nucleus of NGC 4038; he adds that the new object is located on top of a condensed region in the prominent western spiral arm, and eight past images (June 2002–May 2004) show no change in brightness of this region. Gerardy et al. add that a spectrogram (range 400– 820 nm), obtained on Dec. 16.41 with the 2.7-m Harlan J. Smith telescope (+ Imaging Grism Instrument) by A. Bauer, suggests that SN 2004gu is of type Ia; the spectrum roughly resembles that of SN 1999ee slightly before maximum (Hamuy et al. 2002, A.J. 124, 417).

SUPERNOVAE 2004gj AND 2004go Further to IAUC 8453, A. V. Filippenko and R. J. Foley add that their

Dec. 12 spectra show that SN 2004gj (IAUC 8445) is probably of type IIb, perhaps 2 months after the explosion; besides the He I lines, a relatively prominent absorption line at 630 nm is reasonably identified as Hα.

H. Navasardyan, M. Turatto, A. Harutunyan, S. Benetti, and N. Elias-Rosa, Osservatorio Astronomico di Padova; and A. Pastorello, Max-Planck-Institut, report that a spectrogram of SN 2004go (cf. IAUC 8448, 8450), obtained by R. Viotti and C. Rossi on Dec. 7.94 UT with the Asiago 1.8-m telescope (+ AFOSC; range 355–780 nm, resolution 2.4 nm), shows it to be a type-Ia supernova, 3–4 weeks past maximum. The spectrum closely resembles that of SN 2002bo (Benetti et al. 2004, MNRAS 348, 261) at comparable phase.

2004 December 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green